if a,b,c,d are different positive real numbers in H.P., then prove that ad>bc

Grade:12

2 Answers

Riddhish Bhalodia
8 years ago
As the numbers are in HP, their inverse will be in AP and hence
1/b = 1/a + k
1/c = 1/a + 2k
1/d = 1/a + 3k
where k is the constant difference
 


hence
ad>bc
Proved!
arun
8 years ago
i don’t understand the answer given by you actualy it don’t make any sense as you simply said 1/b=1/a+k …...
then you directly wrote ad > bc.